# Computing algebraic Belyi functions on Bring’s curve

### v2

@@ 2.2. An algebraic Belyi function on $X$
Let $X_{0}(50)^{{\rm alg}}$ be the hyperelliptic curve over $\mathbb{Q}$ defined by
@@
Put $T:=y/z=f_{1}/f_{2}$ and $S:=z/w=f_{3}/f_{4}$ so that $T,S\in\mathbb{C}(X_{0}(50))$. By using Gröebner basis for the ideal generated by $zT-y,wS-z$ and (2.2) in $\mathbb{Q}[x,y,z,w,T,S]$, we can eliminate $x,y,z,w$ to obtain a relation $01+S^{2}+2S^{3}-4T-4S^{2}T-T^{2}-S^{2}T^{2}+2S^{3}T^{2}=$ 0. The curve defined by this equation is isomorphic over $\mathbb{Q}$ to (2.4) by
@@
$\beta^{{\rm alg}}_{\Gamma_{B}}:X\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}},\ [x:y:z:w]\mapsto h\Bigg{(}\frac{z+w}{z-w},\frac{2(w^{3}x-2w^{3}y+wxz^{2}-2wyz^{2}+2xz^{3})}{y(z-w)^{3}}\Bigg{)}$
is an algebraic Belyi function defined over $\mathbb{Q}$.
@@
The Belyi function $X(G_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C})$ in question is a non-Galois covering of $\mathbb{P}^{1}(\mathbb{C})$. Thus, it gives rise to a non-regular dessin on $X(G_{B})$.
@@ 2.3. A decomposition of the Jacobian variety over a number field
In this section, we study a number field $L$ such that the Jacobian variety ${\rm Jac}(X)$ of $X$ decomposes completely over $L$. For two abelian varieties $A_{1},A_{2}$ over a number field $K$ and a field extension $M/K$, we write $A_{1}\stackrel{{\scriptstyle M}}{{\sim}}A_{2}$ if $A_{1}$ is isogenous over $M$ to $A_{2}$. For each newform $f$ in $S_{2}(\Gamma_{B})$, we denote by $A_{f}$ the Shimura’s abelian variety over $\mathbb{Q}$ of dimension $[\mathbb{Q}_{f}:\mathbb{Q}]$ where $\mathbb{Q}_{f}$ is the Hecke field of $f$ (cf. [16, Section 3]). We note that $A_{f}$ is $\mathbb{Q}$-simple.
@@
${\rm Jac}(X)\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times J_{0}(50)_{\mathbb{Q}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times E_{{\rm 50.a3}}\times E_{{\rm 50.b3}}\stackrel{{\scriptstyle L}}{{\sim}}E^{4}$
@@
The decomposition over $\mathbb{Q}$ follows from [15, Proposition (2.3)]. It follows from [13] that $A_{f_{{\rm 50.2.b.a}}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}{\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}(E_{1})$ where $E_{1}=E_{100.1-{\rm a}3}:y^{2}+(\varepsilon+1)xy+(\varepsilon+1)y=x^{3}+\varepsilon x^{2},\ \varepsilon:=\displaystyle\frac{1+\sqrt{5}}{2}$ with the $j$-invariant $-\frac{25}{2}$. Here ${\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}$ stands for the Weil-restriction from $\mathbb{Q}(\sqrt{5})$ to $\mathbb{Q}$. Since $E_{1}$ and $E$ are isomorphic each other over $\overline{\mathbb{Q}}$, by [22, Chapter III, Proposition 3.1], there exist $u,r,s,t\in\overline{\mathbb{Q}}$ such that $E\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}E_{1},\ (x,y)\mapsto(u^{2}x+r,u^{3}y+u^{2}sx+r)$. Solving the system of the algebraic equations, we have
@@ 4. A relation between two algebraic models
In this section, we study the field of definition of isomorphisms between $B$, $X$ and the Hulek-Craig curve. It is known that Bring’s curve $B$ is isomorphic to the algebraic model $X$ of $X(\Gamma_{B})$ over $\overline{\mathbb{Q}}$. We would like to discuss a number field $K$ such that $B$ is isomorphic over $K$ to $X$. Obviously, $K$ cannot be $\mathbb{Q}$ since $B(\mathbb{Q})=\emptyset$ while $X(\mathbb{Q})\neq\emptyset$. For another reason, it is known that ${\rm Jac}(B)\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}E^{4}$ (cf. [18, p.87, Exercise-b)]) where $E$ is given in Proposition 2.4. Thus, the field $K$ cannot be $\mathbb{Q}$ since ${\rm Jac}(B)$ is not isogenous over $\mathbb{Q}$ to ${\rm Jac}(X)$ (see also Remark 2.5). We should note that ${\rm Jac}(X)$ is isogenous over $L$ to ${\rm Jac}(B)$, but it is not necessarily true that $X$ is isomorphic over $L$ to $B$. Thus, we need more careful study for our purpose.
@@
Put $w_{i}:=F_{Y}(X,Y,1)\displaystyle\frac{v_{i}}{dX}$ for $i=1,2,3,4$ and let us regard $[f_{1}dq:f_{2}dq:f_{3}dq:f_{4}dq]$ with the coordinates $[x:y:z:w]$ of $X$ in view of the theory of canonical curves. Then, $[x:y:z:w]=[w_{1}:w_{2}:w_{3}:w_{4}]A^{-1}$. Using this, we have a birational map $\psi$ from the singular model of $X^{{\rm HC}}$ defined by $F(x,y,z)$ to $X$ given by $\psi([x_{1}:y_{1}:z_{1}])=$
@@
Then, it yields an isomorphism over $\mathbb{Q}$ from $X^{{\rm HC}}$ to $X$. This result is not obvious from the construction explained in [9, p.82-83]. Thus, we have checked the following result:
@@
Recall that the projective smooth model $X$ over $\mathbb{Q}$ of the modular curve $X(\Gamma_{B})$ from (2.2). It holds that
@@
    Hulek-Craig’s curve $X^{{\rm HC}}$ is isomorphic over $\mathbb{Q}$ to $X$ whose isomorphism is given by (4.1). In particular, Hulek-Craig’s curve is isomorphic over $\mathbb{Q}$ to a canonical model of $X(\Gamma_{B})$ (see [11] for canonical models of modular curves).
Thus, $B$ inherits two algebraic Belyi functions over $\mathbb{Q}(\zeta_{5})$ such that one is non-regular (not a Galois covering) while the other is regular (a Galois covering) (see [7] for the definition of regular Belyi functions).

### v3

@@ 2.2. An algebraic Belyi function on $X(\Gamma_{B})^{{\rm alg}}$
Here $\pi_{1}$ and $\pi_{2}$ are natural projections induced by the inclusion $\Gamma_{B}\subset\Gamma_{0}(50)\subset{\rm SL}_{2}(\mathbb{Z})$. Let $X_{0}(50)^{{\rm alg}}$ be the hyperelliptic curve over $\mathbb{Q}$ defined by
@@
Put $T:=y/z=f_{1}/f_{2}$ and $S:=z/w=f_{3}/f_{4}$ so that $T,S\in\mathbb{C}(X_{0}(50))$. By using Mathematica version 12.1, we can compute a Gröebner basis for the ideal generated by $zT-y,wS-z$ and (2.2) in $\mathbb{Q}[x,y,z,w,T,S]$ such that we can find a relation $01+S^{2}+2S^{3}-4T-4S^{2}T-T^{2}-S^{2}T^{2}+2S^{3}T^{2}=$ 0 in that basis. The curve defined by this equation is isomorphic over $\mathbb{Q}$ to (2.4) by
@@
$\beta^{{\rm alg}}_{\Gamma_{B}}:X(\Gamma_{B})^{{\rm alg}}\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}},\ [x:y:z:w]\mapsto\frac{1}{1728}h\Bigg{(}\frac{z+w}{z-w},\frac{2(w^{3}x-2w^{3}y+wxz^{2}-2wyz^{2}+2xz^{3})}{y(z-w)^{3}}\Bigg{)}$
is an algebraic Belyi function defined over $\mathbb{Q}$. It is also an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C}),\tau\mapsto\frac{1}{1728}j(\tau)$.
@@
The Belyi function $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C})$ in question is a non-Galois covering of $\mathbb{P}^{1}(\mathbb{C})$. Thus, it gives rise to a non-regular dessin on $X(\Gamma_{B})$.
@@ 2.3. A decomposition of the Jacobian variety over a number field
In this section, we study a number field $L$ such that the Jacobian variety ${\rm Jac}(X(\Gamma_{B})^{{\rm alg}})$ of $X(\Gamma_{B})^{{\rm alg}}$ decomposes completely over $L$. For two abelian varieties $A_{1},A_{2}$ over a number field $K$ and a field extension $M/K$, we write $A_{1}\stackrel{{\scriptstyle M}}{{\sim}}A_{2}$ if $A_{1}$ is isogenous over $M$ to $A_{2}$. For each newform $f$ in $S_{2}(\Gamma_{B})$, we denote by $A_{f}$ the Shimura’s abelian variety over $\mathbb{Q}$ of dimension $[\mathbb{Q}_{f}:\mathbb{Q}]$ where $\mathbb{Q}_{f}$ is the Hecke field of $f$ (cf. [16, Section 3]). We note that $A_{f}$ is $\mathbb{Q}$-simple.
@@
${\rm Jac}(X(\Gamma_{B})^{{\rm alg}})\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times J_{0}(50)_{\mathbb{Q}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times E_{{\rm 50.a3}}\times E_{{\rm 50.b3}}\stackrel{{\scriptstyle L}}{{\sim}}E^{4}$
@@
We refer to
https://www.lmfdb.org/EllipticCurve/2.2.5.1/100.1/a/3
linked from
https://www.lmfdb.org/ModularForm/GL2/Q/holomorphic/50/2/b/a/
for [13]. The decomposition over $\mathbb{Q}$ follows from [15, Proposition (2.3)]. It follows from [13] (see the above websites) that $A_{f_{{\rm 50.2.b.a}}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}{\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}(E_{1})$ where $E_{1}=E_{100.1-{\rm a}3}:y^{2}+(\varepsilon+1)xy+(\varepsilon+1)y=x^{3}+\varepsilon x^{2},\ \varepsilon:=\displaystyle\frac{1+\sqrt{5}}{2}$ with the $j$-invariant $-\frac{25}{2}$. Here ${\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}$ stands for the Weil-restriction from $\mathbb{Q}(\sqrt{5})$ to $\mathbb{Q}$. Since $E_{1}$ and $E$ are isomorphic each other over $\overline{\mathbb{Q}}$, by [22, Chapter III, Proposition 3.1], there exist $u,r,s,t\in\overline{\mathbb{Q}}$ such that $E\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}E_{1},\ (x,y)\mapsto(u^{2}x+r,u^{3}y+u^{2}sx+r)$. Solving the system of the algebraic equations, we have
@@ 4. A relation between two algebraic models
In this section, we study the field of definition of isomorphisms between $B$, $X(\Gamma_{B})^{{\rm alg}}$ and the Hulek-Craig curve. It is known that Bring’s curve $B$ is isomorphic to the algebraic model $X(\Gamma_{B})^{{\rm alg}}$ of $X(\Gamma_{B})$ over $\overline{\mathbb{Q}}$. We would like to discuss a number field $K$ such that $B$ is isomorphic over $K$ to $X(\Gamma_{B})^{{\rm alg}}$. Obviously, $K$ cannot be $\mathbb{Q}$ since $B(\mathbb{Q})$ is empty because of the second quadratic equation in (1.1) while obviously $[1:1:1:0]$ is a $\mathbb{Q}$-rational point on $X(\Gamma_{B})^{{\rm alg}}$. For another reason, it is known that ${\rm Jac}(B)\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}E^{4}$ (cf. [18, p.87, Exercise-b)]) where $E$ is given in Proposition 2.4. Thus, the field $K$ cannot be $\mathbb{Q}$ since ${\rm Jac}(B)$ is not isogenous over $\mathbb{Q}$ to ${\rm Jac}(X)$ (see also Remark 2.5). We should note that ${\rm Jac}(X)$ is isogenous over $L$ to ${\rm Jac}(B)$, but it is not necessarily true that $X$ is isomorphic over $L$ to $B$. Thus, we need more careful study for our purpose.
@@
Put $w_{i}:=F_{Y}(X,Y,1)\displaystyle\frac{v_{i}}{dX}$ for $i=1,2,3,4$ and let us regard $[f_{1}dq:f_{2}dq:f_{3}dq:f_{4}dq]$ with the coordinates $[x:y:z:w]$ of $X$ in view of the theory of canonical curves. Then, $[x:y:z:w]=[w_{1}:w_{2}:w_{3}:w_{4}]A^{-1}$. Using this, we have a birational map $\psi$ from the singular model of $X^{{\rm HC}}$ defined by $F(x,y,z)$ to $X(\Gamma_{B})^{{\rm alg}}$ given by $\psi([x_{1}:y_{1}:z_{1}])=$
@@
Then, it yields an isomorphism over $\mathbb{Q}$ from $X^{{\rm HC}}$ to $X(\Gamma_{B})^{{\rm alg}}$. This result is not obvious from the construction explained in [9, p.82-83]. Thus, we have checked the following result:
@@
Recall that the projective smooth model $X(\Gamma_{B})^{{\rm alg}}$ over $\mathbb{Q}$ of the modular curve $X(\Gamma_{B})$ from (2.2). It holds that
@@
    Hulek-Craig’s curve $X^{{\rm HC}}$ is isomorphic over $\mathbb{Q}$ to $X(\Gamma_{B})^{{\rm alg}}$ whose isomorphism is given by (4.1). In particular, Hulek-Craig’s curve is isomorphic over $\mathbb{Q}$ to a canonical model of $X(\Gamma_{B})$ (see [11] for canonical models of modular curves).
Thus, $B$ inherits two algebraic Belyi functions such that one is defined over $\mathbb{Q}(\zeta_{5})$ and non-regular (not a Galois covering) while the other is defined over $\mathbb{Q}$ and regular (a Galois covering) (see [7] for the definition of regular Belyi functions).